.

I. A student calculated the molarity of a solution prepared by dissolving 0.730 mol of table sugar (sucrose, C12H22O11) in 1.8x10^3 mL of water as 4.06x10^-4 M C12H22O11

 


1
Expert's answer
2021-05-20T07:50:31-0400

Solution.

"\nu=0.730mol;"

"V=1.8\sdot10^3mL=1.8L;"

"C=\dfrac{\nu}{V};"

"C=\dfrac{0.730mol}{1.8L}=0.406M;"

The student made a mistake because he did not convert a unit of volume from milliliters to liters. After all, molarity is defined as the number of moles of solute per liter of solution.

Answer: "C=0.406M."